### v1

@@ 0.6 Conclusions and Future Work
This paper started with a general analysis of existing NLP verification approaches, with a view of identifying key components of a general NLP verification methodology, which we distilled as a six component “NLP Verification Pipeline”:
@@
Based on this taxonomy, we formally defined all the components of the pipeline, and implemented it as a tool ANTONIO [21]. ANTONIO allowed us to mix and match different choices for each pipeline component, enabling us to study the effect of various factors and components of the pipeline in a algorithm-independent way. Our main focus was to identify weak or missing parts of the existing NLP verification methodologies. We proposed that NLP verification results must report, in addition to the standard verifiability metric, the following:
@@
  • volumes, generalisability and falsifiability of verified subspaces.
We finished the paper with a study of major pitfalls of NLP verification pipeline from the point of view of LLMs and proposed future wider and more principled use of some standard NLP methods in NLP verification (precision, recall, cosine similarity, ROUGE-N).
Contributions.
The major discoveries of this paper were:
@@
  • In Section 0.4 we showed that NLP verification tasks exhibit a generalisability-verifiability trade-off, and its effects can be very severe, especially if the verification tasks are approached naively. Generalisability is a novel metric we proposed here, and we strongly believe that generalisability should be routinely measured and reported as part of NLP verification pipeline.
@@
  • In Sections 0.4 and 0.5 we showed that it is possible to overcome this trade-off by using several heuristic methods: defining semantic subspaces, training for semantic robustness, choosing the most suited embedding function and filtering with cosine similarity. All of these methods boiled down to defining more precise verifiable subspaces; and all of them can be effectively implemented as part of NLP verification pipelines in the future.
@@
  • In Section 0.5 we demonstrated that accuracy of LLMs, used as embedding functions and generators of perturbations in NLP verification, cannot be taken for granted. They may fail to correctly embed sentences, and they may fail to preserve the semantic meaning of perturbations. Both of these factors influence practical applications of the NLP verification pipeline.
@@
  • In Section 0.5 we demonstrated that even verified subspaces can be semantically falsified: this effect is due to the tension between verification methods that are essentially geometric and the intuitively understood semantic meaning of sentences. By defining the falsifiability metric and using it in our experiments, we found out that the effects of falsifiability do not seem to be severe in practice; but this may vary from one scenario to another. It is important that NLP verification papers are aware of this pitfall, and report faslsifiability alongside verifiability and generalisability.
@@
Finally, we claim as a contribution, a novel coherent methodological framework that allows to include a broad spectrum of NLP, geometric, machine learning, and verification methods under a single umbrella. No previous publication in this domain covered this range and, as this paper illustrated throughout all sections, consistently covering this broad range of methods is crucial for the development of this field.
Future Work.
@@
Following from our in depth analysis of the NLP perspective, we note that even if one has a satisfactory solution to all the issues discussed, there is still the problem of scalability of the available verification algorithms. For example, the most performant neural network verifier, $\alpha\beta$-Crown [118], can only handle networks in the range of tens of millions of trainable parameters. In contrast, in NLP systems, the base model of BERT [32] has around 110 million trainable parameters (considered small compared to modern LLMs – with trainable parameters in the billions!). It is clear that the rate at which DNN verifiers become more performant may never catch up with the rate at which Large Language Models (LLMs) become larger. Then the question arises: how can this pipeline be implemented in the real world?
@@
For future work, we propose to tackle this based on the idea of verifying a smaller DNN (classifier), manageable by verifiers, that can be placed upstream of a complex NLP system as a safeguard. We call this a filter (as mentioned in Section 0.3 and illustrated in Figure 2), and Figure 7 shows how a semantically informed verified filter can be prepended to an NLP system (here, an LLM) to check that safety-critical queries are handled responsibly, e.g. by redirecting the query to a tightly controlled rule-based system instead of a stochastic LLM.
@@
We note that the NLP community has recently proposed guardrails, in order to control the output of LLMs and create safer systems (such as from Open AI, NVIDIA and so on). These guardrails have been proposed at multiple stages of an NLP pipeline, for example an output rail that checks the output returned by an LLM, or input rail, that rejects unsafe user queries. In the figure, we show an application of our filter applied to the user input, and thus creating guarantees that safety critical queries are handled responsibly. In theory these verification techniques we propose may be applied to guardrails at different stages in the system, and we plan to explore this in future work.
A second future direction is to use this work to create NLP verification benchmarks. In 2020, the International Verification of Neural Networks Competition [17] (VNN-COMP) was established to facilitate comparison between existing approaches, bring researchers working on the DNN verification problem together, and help shape future directions of the field. However, on its fourth edition, the competition still lacked NLP verification benchmarks [16]. We propose to use this work for creating NLP verification benchmarks for future editions, to spread the awareness and attention to this field.
@@ Acknowledgements
Authors acknowledge support of EPSRC grant AISEC EP/T026952/1 and NCSC grant Neural Network Verification: in search of the missing spec, Vinay Krupakaran for his help in the evaluating the perturbations and Daniel Kienitz for his help with the eigenspace rotation.

### v3

@@ 6 Conclusions and Future Work
Summary. This paper started with a general analysis of existing NLP verification approaches, with a view of identifying key components of a general NLP verification methodology. We then distilled these into a “NLP Verification Pipeline” consisting of the following six components:
  1. 1.
@@
Based on this taxonomy, we make concrete selections for each component, and implement the pipeline using the tool ANTONIO [16]. ANTONIO allowed us to mix and match different choices for each pipeline component, enabling us to study the effects of varying the components of the pipeline in a algorithm-independent way. Our main focus was to identify weak or missing parts of the existing NLP verification methodologies. We proposed that NLP verification results should report, in addition to the standard verifiability metric, the following:
@@
  • volumes, generalisability and embedding error of verified subspaces.
We finished the paper with a study of the current limitations of the NLP components of the pipeline and proposed possible improvements such as introducing a perturbations filter stage using cosine similarity. One of the major strengths of the pipeline is that each component can be improved individually.
Contributions. The major discoveries of this paper were:
@@
  • In Section 4 we proposed generalisability as a novel metric, and showed that NLP verification methods exhibit a generalisability-verifiability trade-off. The effects of the trade-off can be severe, especially if the verified subspaces are generated naively (e.g. geometrically). We therefore strongly believe that generalisability should be routinely reported as part of NLP verification pipeline.
@@
  • In Sections 4 and 5 we showed that it is possible to overcome this trade-off by using several heuristic methods: defining semantic subspaces, training for semantic robustness, choosing a suitable embedding function and filtering with cosine similarity. All of these methods result in the definition of more precise verifiable subspaces; and all of them can be practically implemented as part of NLP verification pipelines in the future. This is the main positive result of this paper.
@@
  • In Section 5 we demonstrated that there are two key assumptions underlying the definition of subspaces that cannot be taken for granted. Firstly the LLMs, used as embedding functions, may not map semantically similar sentences to similar vectors in the embedding space. Secondly, our algorithmic methods for generating perturbations, whether by LLMs or otherwise, may not always be semantically-preserving operations. Both of these factors influence practical applications of the NLP verification pipeline.
@@
  • In Section 5 we demonstrated that even verified subspaces can exhibit semantic embedding errors: this effect is due to the tension between verification methods that are essentially geometric and the intuitively understood semantic meaning of sentences. By defining the embedding error metric and using it in our experiments, we demonstrated that the effects of embedding errors do not seem to be severe in practice; but this may vary from one scenario to another. It is important that NLP verification papers are aware of this pitfall, and report embedding error alongside verifiability and generalisability.
@@
Finally, we claim as a contribution, a novel, coherent, methodological framework that allows us to include a broad spectrum of NLP, geometric, machine learning, and verification methods under a single umbrella. As illustrated throughout this paper, no previous publication in this domain covered this range and we believe that covering this broad range of methods is crucial for the development of this field.
Limitations and the Role of the Embedding Gap. In this paper, we have shown the effect of the embedding gap in the NLP verification domain. In Section 4.1, we introduced the generalisability metric (Definition 8) to estimate how well subspaces capture semantically similar yet unseen sentences, providing a quantitative lens to examine this challenge. Sections 4.3 and 4.4 demonstrated that geometric subspaces struggle with the verifiability-generalisability trade-off, while semantic subspaces, constructed using semantic-preserving perturbations, show promise in mitigating the embedding gap by better aligning with data semantics.
[…]
This problem is fundamental for machine learning methods deployed in NLP: they always rely on an “embedding function” that maps sentences into real vectors (on which machine learning algorithms operate). There is an implicit assumption that the embedding function works in a way that semantic similarity of sentences is reflected in geometric proximity of their embeddings. However, as general semantic similarity of sentences is not effectively computable, there is no hope that a perfect embedding function will ever be defined. Fundamentally, this is exactly the reason why machine learning (and not symbolic) approaches to NLP proved to be successful: they operate on the assumption that the embedding function is imperfect. As a consequence, any verification pipeline for NLP must include metrics that measure potential embedding errors. Section 5.6 of this paper is entirely devoted to defining this problem in mathematically precise terms and proposing an effective metric for measuring and reporting the severity and effects of the embedding errors.
This paper aims to quantify and address the embedding gap in the NLP domain. For better quantifying the effect of the embedding gap, we proposed precise metrics such as verifiability, generalisability and embedding error, and showed their interplay. This better understanding of the problem gave rise to our main positive result: the method that empirically reduces the effect of the embedding gap. While our findings mark progress, further research is needed to better align geometric representations with semantic meaning, especially in NLP contexts.
@@
Future Work. Following from our in-depth analysis of the NLP perspective, we note that even if one has a satisfactory solution to all the issues discussed, there is still the problem of scalability of the available verification algorithms. For example, the most performant neural network verifier, $\alpha\beta$-Crown [43], can only handle networks in the range of tens of millions of trainable parameters. In contrast, in NLP systems, the base model of BERT [139] has around 110 million trainable parameters (considered small compared to modern LLMs – with trainable parameters in the billions!). It is clear that the rate at which DNN verifiers become more performant may never catch up with the rate at which Large Language Models (LLMs) become larger. Then the question arises: how can this pipeline be implemented in the real world?
@@
For future work, we propose to tackle this based on the idea of verifying a smaller DNN (classifier), manageable by verifiers, that can be placed upstream of a complex NLP system as a safeguard. We call this a filter (as mentioned in Section 3 and illustrated in Figure 2), and Figure 7 shows how a semantically informed verified filter can be prepended to an NLP system (here, an LLM) to check that safety-critical queries are handled responsibly, e.g. by redirecting the query to a tightly controlled rule-based system instead of a stochastic LLM.
@@
We note that the NLP community has recently proposed guardrails, in order to control the output of LLMs and create safer systems (such as from Open AI, NVIDIA and so on). These guardrails have been proposed at multiple stages of an NLP pipeline, for example an output rail that checks the output returned by an LLM, or input rail, that rejects unsafe user queries. In Figure 7, we show an application of our filter applied to the user input, which thus creates guarantees that a subset of safety critical queries are handled responsibly. In theory these verification techniques we propose may be applied to guardrails at different stages in the system, and we plan to explore this in future work.
A second future direction is to use this work to create NLP verification benchmarks. In 2020, the International Verification of Neural Networks Competition [156] (VNN-COMP) was established to facilitate comparison between existing approaches, bring researchers working on the DNN verification problem together, and help shape future directions of the field. However, for some time, the competition still lacked NLP verification benchmarks [157]. In 2024, we contributed a first NLP benchmark to VNN-COMP, using the methodology of this paper, and the tool ANTONIO [158]. We intend to use this work for creating NLP verification benchmarks for future editions, to spread the awareness and attention to this field.
@@ Acknowledgements
Authors acknowledge support of EPSRC grant AISEC EP/T026952/1 and NCSC grant ‘Neural Network Verification: in search of the missing spec’, Vinay Krupakaran for his help in the manual annotation of the perturbed sentences and Daniel Kienitz for his help with the eigenspace rotation. The first author acknowledges the James Watt Scolarship awarded by the Heriot-Watt University.